# Multi-modal Machine Learning Analysis of X-ray Absorption Near-Edge Spectra and Pair Distribution Functions: Performance and Interpretability towards Experimental Design

### v1

@@ 2.1 Charge state classification results
Next, we look at the feature importance to investigate where in the XANES and PDF spectra information that predicts for the charge state resides. The feature importance was measured via the feature’s contribution to the decrease in Gini impurity [louppe2013understanding], weighted by the fraction of samples that used the feature for splitting, and then averaged across all the trees in the forest. We chose this impurity-based approach over the permutation method, which measures the mean decrease in accuracy when each feature column is shuffled, because our spectral features, especially XANES, are multicollinear. When multiple features are correlated, permuting one barely affects the model accuracy because the model still gets information from other correlated features. Hence, the permutation method tends to underestimate the contribution of the correlated features. Indeed, we found that this method ‘turned off’ important regions of the XANES spectra (see SI for results on Ti). This affected XANES more than PDF because XANES features have higher multicollinearity. All further results use the impurity method.
@@
Fig. 1(b) and (c) show impurity-based feature importance for Ti and Fe, respectively. The plots for Mn and Cu are shown in the SI (Figs. S3 and S4). The top left (i) and right (ii) plots in each pane show the feature importance when using XANES and PDF, respectively, separately as input. To help understand which regions of the spectra are highly predictive we show the XANES and PDF spectra from the respective datasets plotted as grey lines, with the darker dashed line showing the average spectrum. For all metals, the post-edge region of XANES is a strong predictor for oxidation state. For Fe, additionally, the pre-edge region also contains important distinguishing information. This is consistent with the findings in Ref. [wilke2001oxidation], where the pre-edge characteristics were used to differentiate Fe2+ from Fe3+. Analysis done on Fe silicates by Guda et al. highlights the importance of both the white-line peak position and post-edge ”pit” energy for Fe oxidation state prediction [guda_understanding_2021], but only the latter is observed here for our oxide dataset.
@@
While the XANES feature-importance plots show that the predictive information is quite spread out, the most important PDF features concentrate around the nearest-neighbor peak (Fig. 1(b)(ii) and (c)(ii)). For example, the most important PDF features for Fe coincide with the leading and trailing edges of the nearest-neighbor peak. We also see the same thing for Mn and Cu (see SI). This is in agreement with the known correlation between the metal cation charge state and the nearest-neighbor bond length [brown1973empirical, brown1985bond]. A change in charge state results in a shift in the nearest neighbor peak in the PDF, which will produce the largest change in the spectrum at the positions of the steepest slopes on the leading and trailing edge of this sharp peak, and this might explain why these features have the most discriminating power. Specifically for Fe, according to Shannon’s table of ionic radii [shannon1976revised], the Fe-O bond lengths for Fe2+ are between 2.01 and 2.18 Å, whereas Fe3+ and Fe4+ form shorter bonds with lengths in the range 1.89-2.045 Å, and 1.985 Å, respectively. Therefore, we might expect the PDF values in the vicinity of $r=1.9$ and 2.3 Å to be indicative of the Fe charge state, as we see in practice.
@@
All peaks in the PDF will be affected by this change in nearest neighbor bond-length, but the higher-$r$ peaks each individually contribute less information to the prediction, as indicated by their lower feature importance. This is presumably because the effects such as shifts on the higher-$r$ peaks is smaller, the peaks themselves are smaller and there is more peak overlap and more averaging across different structures, all of which reduce the information content of any particular point in the curve. However, a strong takeaway is that the feature importance computed from the Random Forest models provides rich information about where in the PDF spectrum information lies for differentiating charge state.
@@
Despite the XANES managing quite well to predict charge state on its own, it is interesting to explore if adding the PDF data leads to overall better predictions. We find that combining XANES and PDF did not improve the charge state prediction over the XANES-only models for 3/4 metals except Fe, as evidence in the F1 histograms in Fig. 1(a) (hatched plots). Only for Fe is there any increase in performance over XANES alone and the improvement is barely larger than the uncertainties.
@@
Multimodal feature importance plots show that information from XANES dominated the charge state prediction whereas the PDF features contributed very little, as shown in Fig. 1(b)(iii) and (c)(iii) for Ti and Fe, respectively (Mn and Cu plots are given in SI). The first 100 features (left half) correspond to points in the XANES spectrum and the second hundred (right half) correspond to points in the PDF. On average, XANES features account for 80% of the total feature importance with some variability across the four datasets. This is somewhat expected since we know the XANES data is rich in information about charge state of the metal ions [henderson2014x, waychunas_synchrotron_1987-1].
@@
The vertical dotted lines mark the locations of important XANES and PDF features when they were used separately. As discussed above, we know that the PDF nearest-neighbor peak (e.g., see the dotted lines in the Fig. 1(b)(ii) and (c)(ii)) contains rich information about the charge state of the metal ion, but this PDF information is largely ‘turned off’ when XANES data are available. With both XANES and PDF present, the XANES feature importance barely changed for Ti and Cu (see SI), and the new peaks still coincide with the dotted lines.
@@
2.2 Coordination number classification results
@@
We now report the results of applying the same approach to predict coordination number. Our datasets (Ti, Mn, Fe, and Cu) contain structures with metal coordination numbers $CN=4$, 5, and 6, where the steps for computing CN is described in Sec. 4. Since coordination number is a structural quantity, naively, we expect PDF to perform well, but XANES to also contain important information on this quantity. As with the charge state classification, we used F1 scores as our performance metric.
@@
Fig. 2(a) shows weighted mean F1 scores of the XANES (left bar) and PDF (middle bar) models for each of the four metals. Surprisingly, although coordination number is closely related to the height of the PDF nearest-neighbor peak, XANES outperformed PDF for most metals (Ti, Mn, and Cu). Fe is the only exception here where XANES and PDF models achieved the same F1 score of 0.93. Both XANES and PDF models performed very well with F1 scores above the trivial classifier baseline, with XANES scores in the range 0.88 (Cu) to 0.95 (Ti) and PDF scores ranging from 0.80 (Cu) to 0.93 (Fe). Most scores here are higher than the F1 scores for charge state classification (see Fig. 1(a)). This indicates that both XANES and PDF contain rich information about the transition metal’s coordination number.
@@
Figure 2: Coordination number classification results: (a) test F1 scores for XANES, PDF, and XANES+PDF models from left to right on the four datasets: Ti, Mn, Fe, and Cu. Black horizontal lines are F1 scores obtained by a trivial classifier that labels all samples as the majority class. Error bars on the test scores represent one standard deviation. Feature importance plots are shown for (b) Ti and (c) Fe datasets, for (i) XANES features, (ii) PDF features, and (iii) XANES+PDF features. Dotted lines mark the locations of prominent features of XANES- and PDF-only feature importances. To better understand if the important features change in the models trained on XANES+PDF, the same lines are drawn on the combined XANES+PDF importance plot (iii) to highlight the changes.
@@
We also investigated whether some aspect of our PDF featurization led to information loss. Initially, we had normalized all the PDFs in the training and testing sets to the range $[-1,1]$. This rescaling might have caused the PDF models’ underperformance, as the amplitude of the first PDF peak is known to contain information about coordination number and we are applying a different scaling to this peak for each PDF pattern in the dataset. To test this hypothesis, we retrained the models using unscaled PDFs and PDFs normalized to $[0,1]$. Using unscaled PDFs slightly improved the coordination number prediction for all metals but still fell short of XANES performance. We also experimented with different normalization methods on XANES, finding that unscaled XANES performed best. Therefore, we chose to use raw XANES and raw PDFs in all of our models including the charge state prediction discussed previously (Fig. 1) and the results shown in Fig. 2.
@@
The pre-edge region of XANES is a strong predictor for coordination number for 3/4 metals (Ti, Mn, and Cu). However, for Cu (see SI Fig. S6), unlike any other metals we investigated, the post-edge region of XANES did not contain important information on the coordination number. Previous studies that predicted coordination numbers from XANES [carbone_classification_2019, torrisi_random_2020, zheng_random_2020, guda_understanding_2021] have reported similar interpretability trends. These include the already known pre-edge importance, particularly high for Ti [farges_ti_1997] but low for Fe [jackson2005multi], along with the notable edge importance observed for Mn and Cu [torrisi_random_2020, zheng_random_2020]. For Mn (see SI Fig. S5), Torrisi et al. reported strong pre-edge significance [torrisi_random_2020], while our results showed a more evenly distributed importance across the full energy range.
@@
On the other hand, the PDF importance plots look very different for Ti and Fe. While the most distinguishing information lies within the nearest-neighbor peak for Ti, the Fe PDF importance extends all the way to $r\sim 10$. We also observed similar PDF ‘fingerprints’ for Mn, but not in Cu (see SI). This indicates that, at least for Fe and Mn, the nearest-neighbor coordination geometry has a ripple effect on the arrangement of atoms further out, making the PDF a good measure of the metal coordination number. One of the most important PDF features for Fe is at $r=2.1$, aligning closely with the nearest-neighbor peak on the average-PDF plot (dark grey line in Fig. 2(b)(ii)) and consistent with the Fe-O bond length range 1.95-2.18 Å for Fe(VI). The Fe-O bond lengths are 2.03 Å or shorter for Fe(IV) and Fe(V) [shannon1976revised], so we think the PDF value at $r=2.1$ helped distinguish Fe(VI) from the rest. The Mn bond lengths are slightly distinctive [shannon1976revised] for different coordination numbers, but the PDF value near $r=1.7$ might indicate the presence of shorter Mn-O bonds that are feasible for Mn(IV) but not for Mn(V) and Mn(VI), both of which form Mn-O bonds that are 1.86 Å or longer.
@@
Unlike the charge state, where the most important PDF features lie on the leading and trailing edges of the nearest-neighbor peak (Fig. 1(b)(ii) and (c)(ii)), the key PDF feature for predicting coordination number aligns with the peak itself. This suggests that the ML model was extracting the coordination number from the height of the nearest-neighbor peak, consistent with what we know from the definition of the PDF. However, it is less clear why the PDF in the higher-$r$ region contains much more information for Fe and Mn, but not Ti and Cu.
@@
The combined XANES and PDF gave results similar to the the charge state prediction, with the F1 scores showing significant improvement only for Fe but not for the other three metals (hatched bar plots in Fig. 2(a)). The feature importance plots for the combined XANES and PDF input for Ti and Fe are shown in Fig. 2(b)(iii) and (c)(iii), respectively (see SI for Mn and Cu plots). For Ti and Cu, XANES features dominated the prediction, accounting for 78% and 62% of the total importance, respectively. For Ti, the importance of the PDF feature near 2 Å is suppressed when the input includes XANES, where its pre-edge region already contains very strongly distinguishing feature for coordination number [farges_ti_1997]. On the other hand, for Fe, all of the strongest PDF peaks are retained when XANES is present without apparent shifts. This is also true for Mn (see SI), where the PDF importance also contains a series of prominent peaks extending into the higher-$r$ region. For Mn, the PDF showed even higher importance than XANES, contributing 63% of the total importance, yet this yielded a smaller improvement in F1 scores compared to Fe, where XANES and PDF had comparable importance when combined.
@@
2.3 Average nearest-neighbor bond-length regression results
In this section, we present the results for the prediction of mean nearest-neighbor bond length. As with coordination number, bond length is a descriptor for the structure of a material, and therefore we expected the model trained solely on the PDF to perform well. Fig. 3(a) shows the RMSE of the test set expressed as a percentage of the mean nearest-neighbor bond length for each metal. Note that shorter bars in this figure indicate better model performance. The test RMSEs ranged from 0.053 to 0.080 Å (see Table S7 in SI), i.e., percentage errors in the range 2.6-3.9% of the mean bond lengths. The RMSEs indicate overall model performance rather than the precision of individual predictions of bond length. For comparison, models that naively extract the position of the nearest-neighbor PDF peak to estimate the mean bond length produced RMSEs between 14 to 20%, much worse than our random forest models.
@@ 2.3 Average nearest-neighbor bond-length regression results
Figure 3: Bond Length Regression Results: (a) test RMSE scores expressed as a percentage of the mean bond length for XANES, PDF, and XANES+PDF models from left to right on the four datasets: Ti, Mn, Fe, and Cu. Note that for these RMSEs smaller numbers indicate better model performance. Error bars on the test scores represent one standard deviation. Feature importance plots are shown for (b) Ti and (c) Fe datasets, for (i) XANES features, (ii) PDF features, and (iii) XANES+PDF features. Dotted lines mark the locations of prominent features of XANES- and PDF-only feature importances. To better understand if the important features change in the models trained on XANES+PDF, the same lines are drawn on the combined XANES+PDF importance plot (iii) to highlight any differences.
Fig. 3(b) and (c) show the feature importance plots for Ti and Fe, respectively. The plots for Mn and Cu are given in the SI. For XANES, the edge features are very important for Ti, Mn, and Cu. Additionally, post-edge features closer to the EXAFS region contribute a lot for Mn and Fe, but less so for Ti and Cu. Torrisi et al. reported very similar interpretability trends for all metals [torrisi_random_2020], but our Mn plot had an additional post-edge peak at 6650 eV (see SI). Interestingly for Fe, while Guda et al. highlighted the significance of both the edge energy and the post-edge features for bond length prediction [guda_understanding_2021], we did not observe a clear edge importance in Fig. 3(c)(i).
Turning to the PDF, the most important features for all metals coincide with the leading and trailing edges of the nearest-neighbor peak, similar to what we saw in the charge state classification (Fig. 1). This behavior can be rationalized on the basis that changes in bond-length result in shifts of the nearest neighbor peak, with the steep leading and trailing edges of the peak being most sensitive to these shifts. This is directly seen in the feature importance when regressing on bond length directly, as well as in charge state, which is strongly correlated with bond length [brown1973empirical, brown1985bond].
@@
As shown in Fig. 3(b)(iii) and (c)(iii), combining XANES and PDF suppressed the PDF feature importance, with XANES features dominating the prediction for all metals. This mirrors the trends observed in the charge state prediction (Fig. 1), though for bond length, some of the PDF importance around the nearest-neighbor peak remained. This effect is most evident for Cu (see SI), where the combined model produced the greatest reduction in the test error, even though its PDF-only model performed relatively worse compared to other metals. Additionally, the XANES importance remains largely unchanged on adding PDF data for Fe and Cu, though we do observe some broadening in the feature importance plots for Ti and Mn.
@@
Contrary to our initial expectations, random forest models trained on XANES outperformed those trained on PDF for all three prediction targets: charge state, coordination number, and mean nearest-neighbor bond length for all four transition metals studied. Though this was expected for electronic properties (e.g., charge state), we were surprised by the superior performance of XANES for structural properties such as coordination number and bond length. In retrospect, it is clear that the electronic spectrum is strongly correlated with local structure, hence XANES contains rich local structural information. However, it is worth investigating why the PDF didn’t perform so well in these structural tasks.
One possible advantage that XANES has is that it yields information only on the environment of the metal ion whereas the total-PDF contains possibly confounding information about the environment of other ions in the system; it is not site specific. We can assess the importance of this factor on simulated data, as all the model training and testing can be repeated by replacing the total-PDF input (which contains distances from every atom to every other atom) with the differential-PDFs (dPDFs) [egami;b;utbp12], which include only distances from the metal ions to all other ions in the system. Although it is not straightforward to measure dPDFs experimentally, they can be measured in principle [aur1983local, kofalt1986differential, waseda1984novel, waseda2002anomalous, petkovLocalStructureIn02000, petkovApplicationDifferentialResonant2018a] and for our current purposes, they are easy to calculate using Diffpy-CMI (see Methods section).
@@
Figure 4: differential-PDF (dPDF) results on Bond Length Regression: (a) test RMSE scores expressed as a percentage of the mean bond length for XANES, PDF, and dPDF models from left to right on the four datasets: Ti, Mn, Fe, and Cu. Note that for these RMSEs smaller numbers indicate better model performance. The XANES and PDF results from Fig. 3 were reproduced here for easy comparison. Error bars on the test scores represent one standard deviation. (b) Change in RMSEs when training the models on multimodal inputs, combining XANES with PDF (solid) or dPDF (black, striped). Negative changes in RMSEs (for Mn, Fe, and Cu) indicate improvement over the XANES-only models. (c) Feature importance plots for the multimodal inputs, XANES+PDF (dashed line) and XANES+dPDF (solid line) for Fe.

### v2

@@ 2.1 Oxidation state classification
Next, we look at the feature importance to investigate where in the XANES and PDF spectra information that predicts for the oxidation state resides. The feature importance score is a measure of each feature’s contribution to the decrease in Gini impurity [44], weighted by the fraction of samples in the training data that used the feature in decision-making for classification. The importance is then averaged across all the decision trees in the forest. Decision trees partition the input parameter space into sub-regions through a series of decisions. Each decision compares whether a feature value is above or below a threshold, which is selected to minimize the Gini impurity (or another loss function, such as RMSE for regression tasks). To estimate feature importance, input features can be ordered based on how much they contribute to the final prediction across all training samples. Note that feature importance is determined at training time and is considered a property of the model and the training data. In our feature importance plots, the score on the $y$-axis represents the individual importance of each feature, and the scores across the input vector sum up to 1.
@@
We chose this impurity-based approach over the permutation method, which measures the mean decrease in accuracy when each feature column is shuffled, because our spectral features, especially XANES, are correlated with each other. When multiple features are correlated, permuting one barely affects the model accuracy because the model still gets information from other correlated features. Hence, the permutation method tends to underestimate the contribution of the correlated features. Indeed, we found that this method ‘turned off’ important regions of the XANES spectra (see SI for results on Ti). This affected XANES more than PDF because XANES features have higher multicollinearity. All further results use the impurity method.
@@
Fig. 1(b) and (c) show impurity-based feature importance for Ti and Fe, respectively. The plots for Mn and Cu are shown in the SI (Figs. S3 and S4). The top left (i) and right (ii) plots in each pane show the feature importance when using XANES and PDF, respectively, separately as input. To help understand which regions of the spectra are highly predictive we show the XANES and PDF spectra from the respective datasets plotted as gray lines, with the darker gray line showing the average spectrum. For all metals, the post-edge region of XANES is a strong predictor for oxidation state. For Fe, additionally, the pre-edge region shows comparable feature importance to the post-edge region in the oxidation state prediction. This Fe pre-edge importance is consistent with the findings in Ref. [45], where the pre-edge characteristics were used to differentiate Fe2+ from Fe3+. Analysis done on Fe silicates by Guda et al. highlights the importance of both the white-line peak position and post-edge ”pit” energy for Fe oxidation state prediction [28], but only the latter is observed here for our oxide dataset.
@@
While the XANES feature-importance plots show that the predictive information is quite spread out, the most important PDF features concentrate around the nearest-neighbor peak (Fig. 1(b)(ii) and (c)(ii)). For example, the most important PDF features for Fe coincide with the leading and trailing edges of the nearest-neighbor peak. We also see the same thing for Mn and Cu (see SI). This is in agreement with the known correlation between the metal cation oxidation state and the nearest-neighbor bond length [46, 47]. A change in oxidation state results in a shift in the nearest neighbor peak in the PDF, which will produce the largest change in the spectrum at the positions of the steepest slopes on the leading and trailing edge of this sharp peak, and this might explain why these features have the most discriminating power. Specifically for Fe, according to Shannon’s table of ionic radii [48], the Fe-O bond lengths for Fe2+ are between 2.01 and 2.18 Å, whereas Fe3+ and Fe4+ form shorter bonds with lengths in the range 1.89-2.045 Å, and 1.985 Å, respectively. Therefore, we might expect the PDF values in the vicinity of $r=1.9$ and 2.3 Å to be indicative of the Fe oxidation state, as we see in practice.
@@
All peaks in the PDF will be affected by this change in nearest neighbor bond-length, but the higher-$r$ peaks each individually contribute less information to the prediction, as indicated by their lower feature importance. This is presumably because the effects such as shifts on the higher-$r$ peaks is smaller, the peaks themselves are smaller and there is more peak overlap and more averaging across different structures, all of which reduce the information content of any particular point in the curve. However, a strong takeaway is that the feature importance computed from the Random Forest models provides rich information about where in the PDF spectrum information lies for differentiating oxidation state.
@@
Despite the XANES managing quite well to predict oxidation state on its own, it is interesting to explore if adding the PDF data leads to overall better predictions. We find that combining XANES and PDF did not improve the oxidation state prediction over the XANES-only models for 3/4 metals except Fe, as evidence in the F1 histograms in Fig. 1(a) (hatched plots). Only for Fe is there any increase in performance over XANES alone and the improvement is barely larger than the uncertainties.
@@
Multimodal feature importance plots show that information from XANES dominated the oxidation state prediction whereas the PDF features contributed very little, as shown in Fig. 1(b)(iii) and (c)(iii) for Ti and Fe, respectively (Mn and Cu plots are given in SI). The first 100 features (left half) correspond to points in the XANES spectrum and the second hundred (right half) correspond to points in the PDF. On average, XANES features account for 80% of the total feature importance with some variability across the four datasets. This is somewhat expected since we know the XANES data is rich in information about oxidation state of the metal ions [49, 50]. Note that the feature importance scores are normalized such that they sum up to 1 across the input vector, so the scaling changed slightly as the combined XANES+PDF input has 200 features instead of 100.
@@
The vertical dotted lines mark the locations of important XANES and PDF features when they were used separately. These locations correspond to local maxima of feature importance, extracted with the find_peaks function in scipy [51]. As discussed above, we know that the PDF nearest-neighbor peak (e.g., see the dotted lines in the Fig. 1(b)(ii) and (c)(ii)) contains rich information about the oxidation state of the metal ion, but this PDF information is largely ‘turned off’ when XANES data are available. With both XANES and PDF present, the XANES feature importance barely changed for Ti and Cu (see SI), and the new peaks still coincide with the dotted lines.
@@
2.2 Coordination number classification
@@
We now report the results of applying the same approach to predict coordination number. Our datasets (Ti, Mn, Fe, and Cu) contain structures with metal coordination numbers $CN=4$, 5, and 6, where the steps for computing CN is described in Sec. 4. Since coordination number is a structural quantity, and PDF is a structural technique, one might expect PDF to perform well on this task, with XANES to also contain important information. As with the oxidation state classification, we used F1 scores as our performance metric.
@@
Fig. 2(a) shows weighted mean F1 scores of the XANES (left bar) and PDF (middle bar) models for each of the four metals. Surprisingly, although coordination number is closely related to the height of the PDF nearest-neighbor peak, XANES outperformed PDF for most metals (Ti, Mn, and Cu). Fe is the only exception here where XANES and PDF models achieved the same F1 score of 0.93. Both XANES and PDF models performed very well with F1 scores above the trivial classifier baseline, with XANES scores in the range 0.88 (Cu) to 0.95 (Ti) and PDF scores ranging from 0.80 (Cu) to 0.93 (Fe). Most scores here are higher than the F1 scores for oxidation state classification (see Fig. 1(a)). This indicates that both XANES and PDF contain rich information about the transition metal’s coordination number.
@@
Figure 2: Coordination number classification results: (a) test F1 scores for XANES (left bar of each triplet, no pattern), PDF (middle bar with circular pattern), and XANES+PDF models (right bar with striped pattern). Results are shown for the four datasets, from left to right: Ti (orange), Mn (green), Fe (blue), and Cu (magenta). Black horizontal lines indicate F1 scores from a trivial classifier that labels all samples as the modal class. Error bars on the test scores represent one standard deviation. Feature importance plots are shown for (b) Ti and (c) Fe datasets, for (i) XANES features, (ii) PDF features, and (iii) XANES+PDF features. In (i) and (ii), the XANES and PDF spectra are plotted in light gray, with scales defined by the secondary $y-$axis on the right. The dark gray line represents the average spectrum. Vertical dotted lines mark the locations of prominent features from the XANES- and PDF-only importance plots. To highlight changes in feature importance for the combined XANES+PDF models, the same lines are shown in (iii) for comparison.
@@
We also investigated whether some aspect of our PDF featurization led to information loss. Initially, we had normalized all the PDFs in the training and testing sets to the range $[-1,1]$. This rescaling might have caused the PDF models’ underperformance, as the amplitude of the first PDF peak is known to contain information about coordination number and we are applying a different scaling to this peak for each PDF pattern in the dataset. To test this hypothesis, we retrained the models using unscaled PDFs and PDFs normalized to $[0,1]$. Using unscaled PDFs slightly improved the coordination number prediction for all metals but still fell short of XANES performance. We also experimented with different normalization methods on XANES, finding that unscaled XANES performed best. Therefore, we chose to use raw XANES and raw PDFs in all of our models including the oxidation state prediction discussed previously (Fig. 1) and the results shown in Fig. 2.
@@
The pre-edge region of XANES is a strong predictor for coordination number for 3/4 metals (Ti, Mn, and Cu). However, for Cu (see SI Fig. S6), unlike any other metals we investigated, the post-edge region of XANES did not contain important information on the coordination number. Previous studies that predicted coordination numbers from XANES [22, 24, 23, 28] have reported similar interpretability trends. These include the already known pre-edge importance, particularly high for Ti [52] but low for Fe [53], along with the notable edge importance observed for Mn and Cu [24, 23]. For Mn (see SI Fig. S5), Torrisi et al. reported strong pre-edge significance [24], while our results showed a more evenly distributed importance across the full energy range.
@@
On the other hand, the PDF importance plots look very different for Ti and Fe. While the most distinguishing information lies within the nearest-neighbor peak for Ti, the Fe PDF importance extends all the way to $r\sim 10$. We also observed similar PDF ‘fingerprints’ for Mn, but not in Cu (see SI). This indicates that, at least for Fe and Mn, the nearest-neighbor coordination geometry has a ripple effect on the arrangement of atoms further out, making the PDF a good measure of the metal coordination number. One of the most important PDF features for Fe is at $r=2.1$, aligning closely with the nearest-neighbor peak on the average-PDF plot (dark gray line in Fig. 2(b)(ii)) and consistent with the Fe-O bond length range 1.95-2.18 Å for Fe(VI). The Fe-O bond lengths are 2.03 Å or shorter for Fe(IV) and Fe(V) [48], so we think the PDF value at $r=2.1$ helped distinguish Fe(VI) from the rest. The Mn bond lengths are slightly distinctive [48] for different coordination numbers, but the PDF value near $r=1.7$ might indicate the presence of shorter Mn-O bonds that are feasible for Mn(IV) but not for Mn(V) and Mn(VI), both of which form Mn-O bonds that are 1.86 Å or longer.
@@
Unlike the oxidation state, where the most important PDF features lie on the leading and trailing edges of the nearest-neighbor peak (Fig. 1(b)(ii) and (c)(ii)), the key PDF feature for predicting coordination number aligns with the peak itself. This suggests that the ML model was extracting the coordination number from the height of the nearest-neighbor peak, consistent with what we know from the definition of the PDF. However, it is less clear why the PDF in the higher-$r$ region contains much more information for Fe and Mn, but not Ti and Cu.
@@
The combined XANES and PDF gave results similar to the the oxidation state prediction, with the F1 scores showing significant improvement only for Fe but not for the other three metals (hatched bar plots in Fig. 2(a)). The feature importance plots for the combined XANES and PDF input for Ti and Fe are shown in Fig. 2(b)(iii) and (c)(iii), respectively (see SI for Mn and Cu plots). For Ti and Cu, XANES features dominated the prediction, accounting for 78% and 62% of the total importance, respectively. For Ti, the importance of the PDF feature near 2 Å is suppressed when the input includes XANES, where its pre-edge region already contains very strongly distinguishing feature for coordination number [52]. On the other hand, for Fe, all of the strongest PDF peaks are retained when XANES is present without apparent shifts. This is also true for Mn (see SI), where the PDF importance also contains a series of prominent peaks extending into the higher-$r$ region. For Mn, the PDF showed even higher importance than XANES, contributing 63% of the total importance, yet this yielded a smaller improvement in F1 scores compared to Fe, where XANES and PDF had comparable importance when combined.
@@
2.3 Mean nearest-neighbor bond-length regression
In this section, we present the results for the prediction of mean nearest-neighbor bond length. As with coordination number, bond length is a descriptor for the structure of a material where one might expect the model trained solely on the PDF to perform well. Fig. 3(a) shows the RMSE of the test set expressed as a percentage of the mean nearest-neighbor bond length for each metal. Note that shorter bars in this figure indicate better model performance. The test RMSEs ranged from 0.053 to 0.080 Å (see Table S8 in SI), i.e., percentage errors in the range 2.6-3.9% of the mean bond lengths. The RMSEs indicate overall model performance rather than the precision of individual predictions of bond length. For comparison, models that naively extract the position of the nearest-neighbor PDF peak to estimate the mean bond length produced RMSEs between 14 to 20%, much worse than our random forest models.
@@ 2.3 Mean nearest-neighbor bond-length regression
Figure 3: Bond Length Regression Results: (a) test RMSE scores expressed as a percentage of the mean bond length for XANES (left bar of each triplet, no pattern), PDF (middle bar with circular pattern), and XANES+PDF models (right bar with striped pattern). Results are shown for the four datasets, from left to right: Ti (orange), Mn (green), Fe (blue), and Cu (magenta). Note that for these RMSEs smaller numbers indicate better model performance. Error bars on the test scores represent one standard deviation. Feature importance plots are shown for (b) Ti and (c) Fe datasets, for (i) XANES features, (ii) PDF features, and (iii) XANES+PDF features. In (i) and (ii), the XANES and PDF spectra are plotted in light gray, with scales defined by the secondary $y-$axis on the right. The dark gray line represents the average spectrum. Vertical dotted lines mark the locations of prominent features from the XANES- and PDF-only importance plots. To highlight changes in feature importance for the combined XANES+PDF models, the same lines are shown in (iii) for comparison.
Fig. 3(b) and (c) show the feature importance plots for Ti and Fe, respectively. The plots for Mn and Cu are given in the SI. For XANES, the edge features are very important for Ti, Mn, and Cu. Additionally, post-edge features closer to the EXAFS region contribute a lot for Mn and Fe, but less so for Ti and Cu. Torrisi et al. reported very similar interpretability trends for all metals [24], but our Mn plot had an additional post-edge peak at 6650 eV (see SI). Interestingly for Fe, while Guda et al. highlighted the significance of both the edge energy and the post-edge features for bond length prediction [28], we did not observe a clear edge importance in Fig. 3(c)(i).
Turning to the PDF, the most important features for all metals coincide with the leading and trailing edges of the nearest-neighbor peak, similar to what we saw in the oxidation state classification (Fig. 1). This behavior can be rationalized on the basis that changes in bond-length result in shifts of the nearest neighbor peak, with the steep leading and trailing edges of the peak being most sensitive to these shifts. This is directly seen in the feature importance when regressing on bond length directly, as well as in oxidation state, which is strongly correlated with bond length [46, 47].
@@
As shown in Fig. 3(b)(iii) and (c)(iii), combining XANES and PDF suppressed the PDF feature importance, with XANES features dominating the prediction for all metals. This mirrors the trends observed in the oxidation state prediction (Fig. 1), though for bond length, some of the PDF importance around the nearest-neighbor peak remained. This effect is most evident for Cu (see SI), where the combined model produced the greatest reduction in the test error, even though its PDF-only model performed relatively worse compared to other metals. Additionally, the XANES importance remains largely unchanged on adding PDF data for Fe and Cu, though we do observe some broadening in the feature importance plots for Ti and Mn.
@@
Contrary to our initial expectations, random forest models trained on XANES outperformed those trained on PDF for all three prediction targets: oxidation state, coordination number, and mean nearest-neighbor bond length for all four transition metals studied. Though this was expected for electronic properties (e.g., oxidation state), we were surprised by the superior performance of XANES for structural properties such as coordination number and bond length. Because of the intimate relationship between the atomic positions (structural signal) and the resulting electronic spectrum (spectroscopic signal) it is clear that structural techniques can be sensitive to electronic properties, and vice versa. Hence, XANES contains rich local structural information. However, it is worth investigating why the PDF didn’t perform so well in these structural tasks.
One possible advantage that XANES has is that it yields information only on the environment of the metal ion whereas the total-PDF contains possibly confounding information about the environment of other ions in the system; it is not site specific. We can assess the importance of this factor on simulated data, as all the model training and testing can be repeated by replacing the total-PDF input (which contains distances from every atom to every other atom) with the differential-PDFs (dPDFs) [1], which include only distances from the metal ions to all other ions in the system. Although it is not straightforward to measure dPDFs experimentally, they can be measured in principle [54, 55, 56, 57, 58, 59] and for our current purposes, they are easy to calculate using Diffpy-CMI (see Methods section).
@@
Figure 4: differential-PDF (dPDF) results on Bond Length Regression: (a) test RMSE scores expressed as a percentage of the mean bond length for XANES (left bar of each triplet, light color with no pattern), PDF (middle bar with circular pattern), and dPDF models (right bar, black). Results are shown for the four datasets, from left to right: Ti (orange), Mn (green), Fe (blue), and Cu (magenta). Note that for these RMSEs smaller numbers indicate better model performance. The XANES and PDF results from Fig. 3 are reproduced here for easy comparison. Error bars on the test scores represent one standard deviation. (b) Change in RMSEs when training the models on multimodal inputs, combining XANES with PDF (solid) or dPDF (black, striped). Negative changes in RMSEs (for Mn, Fe, and Cu) indicate an improvement over the XANES-only models. (c) Feature importance plots for the multimodal inputs, comparing XANES+PDF (dashed line) and XANES+dPDF (solid line) models for Fe.